# Distributed Solving of Linear Quadratic Optimal Controller with Terminal State Constraint

### v2

@@ II Problem formulation
In this paper, we study a novel case with a sensor network in the system that consists of $N$ agents or nodes, and the communication network between the agents is represented by the undirected topological graph $\mathscr{G}=\{\mathscr{V},\mathscr{E},\mathscr{A}\}$. In particular, each agent $i,i=1,2,\ldots,N$ can only communicate with its neighbors $N_{i}$ without needing global topology information. The graph is assumed to be connected. For more details about the graph refer to [28].
@@
III Preliminaries
III-A Explicit solution by using all the information
@@
Definition 1: A state $x_{T}$ is said to be reachable at time $T$ from the initial state $x_{0}$ of system (1) if there exists a controller $u(t)$, $t\in[0,T]$ such that the solution $x(t)$ of the system (1) satisfies $x(T)=x_{T}$.
@@
$u^{*}(t)=R^{-1}B^{\prime}P(t)x^{*}(t)-R^{-1}B^{\prime}\beta(t),$ (16)
@@
Proof: The proof follows directly from Theorem 1 in [33] and is thus omitted here. $\blacksquare$
III-B Numerical solution by using all the information
@@
Proof: The proof can be directly obtained from Theorem 7.2 in [34] and is thus omitted here. $\blacksquare$
@@
Proof: The result follows directly from Lemma 1. $\blacksquare$
From Lemma 1 and Corollary 1, we have obtained the numerical solution to (13) and (14). Accordingly, the optimal solution (16) can be numerically obtained, and then the detailed algorithm is given in Algorithm 1.
@@ 5. Solve the equation
Proof: From Lemma 1, there always exist an integer $n$ and a sufficient small constant $\varepsilon>0$ such that $\|P^{n}(t)-P(t)\|\leq\varepsilon,t\in[0,T]$. This further implies that there exist sufficient small constants $\varepsilon_{1}>0,\varepsilon_{2}>0$ such that
@@ IV Main results
$Z_{i,k}^{n}(t)$ $=$ $Z_{i,k-1}^{n}(t)+\alpha_{k}[A_{i}-M_{i}M_{i}^{\prime}P_{i,k-1}^{n-1}(t)$ (31)
$-Z_{i,k-1}^{n}(t)]+\sum_{j\in N_{i}}[Z_{j,k-1}^{n}(t)-Z_{i,k-1}^{n}(t)],$
$V_{i,k}^{n}(t)$ $=$ $V_{i,k-1}^{n}(t)+\alpha_{k}[Q_{i}+P_{i,k-1}^{n-1}(t)M_{i}M_{i}^{\prime}P_{i,k-1}^{n-1}(t)$ (32)
$-V_{i,k-1}^{n}(t)]+\sum_{j\in N_{i}}[V_{j,k-1}^{n}(t)-V_{i,k-1}^{n}(t)],$
while $\alpha_{k}$ is a step size satisfying $\sum_{k=0}^{\infty}\alpha_{k}=\infty$, $\sum_{k=0}^{\infty}\alpha_{k}^{2}\leq\infty$, and $Z_{i,0}^{n}(t)=0$, $V_{i,0}^{n}(t)=0$.
@@
Proof: The detailed proof is given in Appendix A. $\blacksquare$
@@
Proof: The result follows directly from Lemma 2. $\blacksquare$
@@
Proof: The proof is similar to Lemma 2 and is omitted here. $\blacksquare$
@@
$-\lambda_{i,k-1}^{*}\}+\sum_{j\in N_{i}}[\lambda_{j,k-1}^{*}-\lambda_{i,k-1}^{*}].$
@@
Proof: The proof is similar to Lemma 2 and is omitted here. $\blacksquare$
@@
Proof: The proof is similar to Lemma 2 and is omitted here. $\blacksquare$
From Lemmas 2-5 and Corollary 2, we have established the distributed solutions to (13) and (14). This enables the distributed design of the optimal controller (16) as shown in Algorithm 2.
@@
    $+\sum_{j\in N_{i}}[W_{j,\varpi-1}-W_{i,\varpi-1}].$
@@
    $-\lambda_{i,q-1}^{*}\}+\sum_{j\in N_{i}}[\lambda_{j,q-1}^{*}-\lambda_{i,q-1}^{*}].$
@@
Theorem 3
$\|x_{i,w}^{*}(t)-x^{*}(t)\|\leq\psi_{1},$ (56)
$\|u_{i}^{*}(t)-u^{*}(t)\|\leq\psi_{2},$ (57)
Proof. From Lemmas 1-2 and Corollary 2, there always exists integer $n$, $k$ and a sufficient small constant $\sigma_{1}>0$ such that $\|P_{i,k}^{n}(t)-P(t)\|\leq\sigma_{1},t\in[0,T]$. Together with Lemmas 3-4, this further implies that there exist iteration numbers $\varpi,q$ and sufficient small constants $\sigma_{i}>0,i=2,\cdots,4$ such that
$\|W_{i,\varpi}-W\|\leq\sigma_{2},$
$\|\lambda_{i,q}^{*}-\lambda^{*}\|\leq\sigma_{3},$
$\|\beta_{i}(t)-\beta(t)\|\leq\sigma_{4},$
Accordingly, by using Lemma 5, there exist iteration numbers $w$ and sufficient small constants $\psi_{1}>0$, $\psi_{2}>0$ such that the outputs $x_{i,w}^{*}(t)$ and $u_{i}^{*}(t)$ obtained by running Algorithm 2 satisfies (56) and (57). This completes the proof. $\blacksquare$
In this case, the optimal controller $u^{\diamond}(t)$ is
$u^{\diamond}(t)=-R^{-1}B^{\prime}P(t)x^{\diamond}(t),$ (58)
@@
    $+\sum_{j\in N_{i}}[W_{j,\bar{\varpi}-1}-W_{i,\bar{\varpi}-1}].$
@@
    $+\sum_{j\in N_{i}}[x_{j,\bar{w}-1}^{\diamond}(t)-x_{i,\bar{w}-1}^{\diamond}(t)].$
@@ V Simulation Example
Next, by running Algorithm 2 with step size $\alpha_{k}=\frac{1}{k}$, we obtain $P_{i,k}^{n}(t)$, $W_{i,\varpi}$, $\lambda_{i,q}^{*}$, $\beta_{i}(t)$, $x_{i,w}^{*}(t)$ and $u_{i}^{*}(t)$, and the corresponding trajectories are shown in Figure 3-Figure 8. It is obvious that all the agents achieve consensus and converge to the centralized solutions from Figure 7-Figure 8. In particular, the trajectories of the agent achieve the predefined target formation at the given time $T=2$s as shown in Figure 9.
@@
By running Algorithm 2 with $\alpha_{k}=\frac{1}{k}$, we obtain $P_{i,k}^{n}(t)$, $W_{i,\varpi}$, $\lambda_{i,q}^{*}$, $\beta_{i}(t)$, $x_{i,w}^{*}(t)$ and $u_{i}^{*}(t)$, and the corresponding trajectories are shown in Figure 11-Figure 16. It is obvious that the distributed solutions converge to the centralized solutions Figure 15-Figure 16.
@@ VI Conclusions
where $\mathcal{A}=I_{N}-\mathcal{L}$, $\mathcal{L}$ is the Laplacian matrix [31]. It is obvious that $\mathcal{Z}_{k}^{1}(t)$ and $\mathcal{V}_{k}^{1}(t)$ are bounded. In fact, by defining $\bar{\mathcal{Z}}_{k}^{1}(t)=(U^{\prime}\otimes I_{n})\mathcal{Z}_{k}^{1}(t)$ and $\bar{\mathcal{V}}_{k}^{1}(t)=(U^{\prime}\otimes I_{n})\mathcal{V}_{k}^{1}(t)$, where $U$ is orthogonal matrix such that $U^{\prime}\mathcal{L}U=diag\{\lambda_{1},\cdots,\lambda_{N}\}$ and $0=\lambda_{1}\leq\lambda_{2}\leq\cdots\leq\lambda_{N}$ are the eigenvalues of $\mathcal{L}$. According to (86)-(87), we have that
$\bar{\mathcal{Z}}_{k}^{1}(t)$ $=$ $[(I-\alpha_{k}I-U^{\prime}\mathcal{L}U)\otimes I_{n}]\bar{\mathcal{Z}}_{k-1}^{1}(t)$
@@
$\bar{\mathcal{V}}_{k}^{1}(t)$ $=$ $[(I-\alpha_{k}I-U^{\prime}\mathcal{L}U)\otimes I_{n}]\bar{\mathcal{V}}_{k-1}^{1}(t)$
@@
In view of the fact that the graph is connected, it yields that

### v3

@@ 2 Problem formulation
In this paper, we study a novel case with a sensor network in the system that consists of $N$ agents or nodes, and the communication network between the agents is represented by the undirected topological graph $\mathscr{G}=\{\mathscr{V},\mathscr{E},\mathscr{A}\}$. In particular, each agent $i,i=1,2,\ldots,N$ can only communicate with its neighbors $N_{i}$. The graph is assumed to be connected. For more details about the graph refer to [28].
@@
3 Preliminaries
3.1 Explicit solution by using all the information
@@
Definition 1.
A state $x_{T}$ is said to be reachable at time $T$ from the initial state $x_{0}$ of system (1) if there exists a controller $u(t)$, $t\in[0,T]$ such that the solution $x(t)$ of the system (1) satisfies $x(T)=x_{T}$.
@@
$u^{*}(t)=-R^{-1}B^{\prime}P(t)x^{*}(t)-R^{-1}B^{\prime}\beta(t),$ (16)
@@
Proof. The proof follows directly from Theorem 1 in [33] and is thus omitted here. $\blacksquare$
3.2 Numerical solution by using all the information
@@
Proof. The proof can be directly obtained from Theorem 7.2 in [34] and is thus omitted here. $\blacksquare$
@@
Proof. The result follows directly from Lemma 3.1. $\blacksquare$
From Lemma 3.1 and Corollary 3.2, we have obtained the numerical solution to (13) and (14). Accordingly, the optimal solution (16) can be numerically obtained, and then the detailed algorithm is given in Algorithm 1.
@@ 5. Solve the equation
Proof. From Lemma 3.1, there always exist an integer $n$ and a sufficient small constant $\varepsilon>0$ such that $\|P^{n}(t)-P(t)\|\leq\varepsilon,t\in[0,T]$. This further implies that there exist sufficient small constants $\varepsilon_{1}>0,\varepsilon_{2}>0$ such that
@@ 4 Main results
$Z_{i,k}^{n}(t)=$ $Z_{i,k-1}^{n}(t)+\alpha_{k}[A_{i}-M_{i}M_{i}^{\prime}P_{i,k-1}^{n-1}(t)-Z_{i,k-1}^{n}(t)]$
$+\frac{1}{\gamma}\sum_{j\in N_{i}}[Z_{j,k-1}^{n}(t)-Z_{i,k-1}^{n}(t)],$ (31)
$V_{i,k}^{n}(t)=$ $V_{i,k-1}^{n}(t)+\alpha_{k}[Q_{i}+P_{i,k-1}^{n-1}(t)M_{i}M_{i}^{\prime}P_{i,k-1}^{n-1}(t)$
$-V_{i,k-1}^{n}(t)]+\frac{1}{\gamma}\sum_{j\in N_{i}}[V_{j,k-1}^{n}(t)$
$-V_{i,k-1}^{n}(t)],$ (32)
while $\alpha_{k}$ is a step size satisfying $\sum_{k=0}^{\infty}\alpha_{k}=\infty$, $\sum_{k=0}^{\infty}\alpha_{k}^{2}\leq\infty$, $\gamma>0$, and $Z_{i,0}^{n}(t)=0$, $V_{i,0}^{n}(t)=0$.
@@
Proof. The detailed proof is given in Appendix A. $\blacksquare$
@@
Proof. The result follows directly from Lemma 4.1. $\blacksquare$
@@
Proof. The proof is similar to Lemma 2 and is omitted here. $\blacksquare$
@@
$-\lambda_{i,k-1}^{*}\}+\frac{1}{\gamma}\sum_{j\in N_{i}}[\lambda_{j,k-1}^{*}-\lambda_{i,k-1}^{*}].$
@@
Proof. The proof is similar to Lemma 2 and is omitted here. $\blacksquare$
@@
Proof. The proof is similar to Lemma 2 and is omitted here. $\blacksquare$
From Lemmas 4.1-4.5 and Corollary 4.2, we have established the distributed solutions to (13) and (14). This enables the distributed design of the optimal controller (16) as shown in Algorithm 2.
Theorem 3.
$\|x_{i,w}^{*}(t)-x^{*}(t)\|\leq\psi_{1},$ (56)
$\|u_{i}^{*}(t)-u^{*}(t)\|\leq\psi_{2},$ (57)
Proof. From Lemmas 3.1-4.1 and Corollary 4.2, there always exists integer $n$, $k$ and a sufficient small constant $\sigma_{1}>0$ such that $\|P_{i,k}^{n}(t)-P(t)\|\leq\sigma_{1},t\in[0,T]$. Together with Lemmas 4.3-4.4, this further implies that there exist iteration numbers $\varpi,q$ and sufficient small constants $\sigma_{i}>0,i=2,\cdots,4$ such that
$\|W_{i,\varpi}-W\|\leq\sigma_{2},$
$\|\lambda_{i,q}^{*}-\lambda^{*}\|\leq\sigma_{3},$
$\|\beta_{i}(t)-\beta(t)\|\leq\sigma_{4},$
Accordingly, by using Lemma 4.5, there exist iteration numbers $w$ and sufficient small constants $\psi_{1}>0$, $\psi_{2}>0$ such that the outputs $x_{i,w}^{*}(t)$ and $u_{i}^{*}(t)$ obtained by running Algorithm 2 satisfies (56) and (57). This completes the proof. $\blacksquare$
In this case, the optimal controller $u^{\diamond}(t)$ is
$u^{\diamond}(t)=-R^{-1}B^{\prime}P(t)x^{\diamond}(t),$ (58)
@@
    $+\frac{1}{\gamma}\sum_{j\in N_{i}}[W_{j,\varpi-1}-W_{i,\varpi-1}].$
@@
    $-\lambda_{i,q-1}^{*}\}+\frac{1}{\gamma}\sum_{j\in N_{i}}[\lambda_{j,q-1}^{*}-\lambda_{i,q-1}^{*}].$
@@
    $+\frac{1}{\gamma}\sum_{j\in N_{i}}[W_{j,\bar{\varpi}-1}-W_{i,\bar{\varpi}-1}].$
@@
    $+\frac{1}{\gamma}\sum_{j\in N_{i}}[x_{j,\bar{w}-1}^{\diamond}(t)-x_{i,\bar{w}-1}^{\diamond}(t)].$
@@ 5 Simulation Example
Next, by running Algorithm 2 with step size $\alpha_{k}=\frac{1}{k}$ and $\gamma=0.4$, we obtain $P_{i,k}^{n}(t)$, $W_{i,\varpi}$, $\lambda_{i,q}^{*}$, $\beta_{i}(t)$, $x_{i,w}^{*}(t)$ and $u_{i}^{*}(t)$, and the corresponding trajectories are shown in Figure 3-Figure 8. It is obvious that all the agents achieve consensus and converge to the centralized solutions from Figure 7-Figure 8. In particular, the trajectories of the agent achieve the predefined target formation at the given time $T=2$s as shown in Figure 9.
@@
By running Algorithm 2 with $\alpha_{k}=\frac{1}{k}$, $\gamma=0.4$, we obtain $P_{i,k}^{n}(t)$, $W_{i,\varpi}$, $\lambda_{i,q}^{*}$, $\beta_{i}(t)$, $x_{i,w}^{*}(t)$ and $u_{i}^{*}(t)$, and the corresponding trajectories are shown in Figure 11-Figure 16. It is obvious that the distributed solutions converge to the centralized solutions Figure 15-Figure 16.
@@ 6 Conclusions
where $\mathcal{A}=I_{N}-\frac{1}{\gamma}\mathcal{L}$, $\mathcal{L}$ is the Laplacian matrix [31]. It is obvious that $\mathcal{Z}_{k}^{1}(t)$ and $\mathcal{V}_{k}^{1}(t)$ are bounded. In fact, by defining $\bar{\mathcal{Z}}_{k}^{1}(t)=(U^{\prime}\otimes I_{n})\mathcal{Z}_{k}^{1}(t)$ and $\bar{\mathcal{V}}_{k}^{1}(t)=(U^{\prime}\otimes I_{n})\mathcal{V}_{k}^{1}(t)$, where $U$ is orthogonal matrix such that $U^{\prime}\mathcal{L}U=diag\{\lambda_{1},\cdots,\lambda_{N}\}$ and $0=\lambda_{1}\leq\lambda_{2}\leq\cdots\leq\lambda_{N}$ are the eigenvalues of $\mathcal{L}$. According to (86)-(87), we have that
$\bar{\mathcal{Z}}_{k}^{1}(t)$ $=$ $[(I-\frac{1}{\gamma}U^{\prime}\mathcal{L}U)\otimes I_{n}]\bar{\mathcal{Z}}_{k-1}^{1}(t)$
@@
$\bar{\mathcal{V}}_{k}^{1}(t)$ $=$ $[(I-\frac{1}{\gamma}U^{\prime}\mathcal{L}U)\otimes I_{n}]\bar{\mathcal{V}}_{k-1}^{1}(t)$
@@
Since the graph is connected, there exists a constant $\gamma>0$ such that the following inequality holds